Pheromones are compounds secreted by females of many insect species to attract mates. Typically, 1.00 × 10−8 g of a pheromone is sufficient to reach all targeted males within a radius of 0.500 mi. Calculate the density of the pheromone (in grams per liter) in a cylindrical air space having a radius of 0.500 mi and a height of 39.0 ft (volume of a cylinder of radius r and height h is πr2h). Enter your answer in scientific notation.

Answer: ___x10^__ g/L
1
Expert's answer
2019-10-18T08:39:49-0400

V= pr^2 x 2h= 3.14 x 8046,72^2 x 2 x 1.642872 = 4.81622366e-5 dm3 = 4.81622366e-5 L.

p= m/V = 1.00 × 10−8/4.81622366e-5= 0.000207631553 g/L.